OpenStudy (anonymous):

Which of the following is the equation of a line with a slope of -5 and passing through (4,-2) A.Y-2=-5(x+4) B.Y+2=-5(x+4) C.Y+2=-5(x+4) D.Y-2=-5(x+4)

OpenStudy (anonymous):

I hate graphing like questions..so I'm going to show up a lot..well I don't understand them very well. All I understood was "-5-"....

OpenStudy (katrinakaif):

Your slope would substitute to be: "m" in the formula: (y -y1) = m (x-x1) Your given points: (4, -2) is the substitution for (x1,y1) Thus subsituting: (y-(-2) = -5(x -4) t + 2 = -5 (x -4)

OpenStudy (katrinakaif):

All your options are invalid.

Directrix (directrix):

Are B and C the same options? My eyes may be failing me.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

A and D are also the same, the packet was typed up like that.

Directrix (directrix):

Check the question. It is so easy to do a typo.

OpenStudy (anonymous):

your options are incorrect.........check again